OpenStudy (anonymous):

The distance formula can be used to prove a quadrilateral has? a right angle congruent angles parallel sides congruent sides

jimthompson5910 (jim_thompson5910):

hint: The distance formula can be used to find the lengths of the sides.

jimthompson5910 (jim_thompson5910):

Since finding the distance from point X to point Y is the same as finding the length of segment XY

OpenStudy (anonymous):

so basicly its used for congruent sides?

OpenStudy (anonymous):

so its congruent sides?

jimthompson5910 (jim_thompson5910):

correct, if you find that all the sides are the same length, then they are congruent

jimthompson5910 (jim_thompson5910):

if all 4 sides are congruent, then you have a rhombus

OpenStudy (anonymous):

thank you very much

jimthompson5910 (jim_thompson5910):

if you have one pair of congruent sides, and then another pair of congruent sides (not congruent to the first two), then you have a parallelogram

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Line AB contains points A (−1, 3) and B (5, 3). Line AB has a slope that is zero undefined positive negative i have a feeling this answer is zero lol

jimthompson5910 (jim_thompson5910):

what gives you this feeling?

OpenStudy (anonymous):

because im looking a the y's and there the same cause i think they lie on the same line

jimthompson5910 (jim_thompson5910):

and that line is horizontal all horizontal lines have a slope of 0 good

jimthompson5910 (jim_thompson5910):

you can also use the slope formula \[\Large m = \frac{y_2-y_1}{x_2-x_1}\] and you'll also find it's 0

OpenStudy (anonymous):

yes exactly thats how i got it also

OpenStudy (anonymous):

The equation of line GH is y = 4x − 3. Write an equation of a line perpendicular to line GH in slope-intercept form that contains point (2, 3). y = 4x − 3 y = −4x − 5 11/4 x − 3 y =-1/4 x + 7/2

OpenStudy (anonymous):

can you help me on this

jimthompson5910 (jim_thompson5910):

How far did you get?

jimthompson5910 (jim_thompson5910):

What's the slope of the given line?

OpenStudy (anonymous):

y=4x-3?

jimthompson5910 (jim_thompson5910):

that's the given equation of the given line

jimthompson5910 (jim_thompson5910):

what's the slope of that line?

OpenStudy (anonymous):

idk help me please

jimthompson5910 (jim_thompson5910):

hint: y = mx+b has a slope of m

jimthompson5910 (jim_thompson5910):

So something like y = 7x - 10 has a slope of 7 (since m = 7)

OpenStudy (anonymous):

then it would be 4? that would be slope

jimthompson5910 (jim_thompson5910):

good, which is the same as 4/1 Now flip the fraction to get 1/4 Then flip the sign to get -1/4 So that's the perpendicular slope

jimthompson5910 (jim_thompson5910):

Given Slope: 4 Perpendicular Slope: -1/4

OpenStudy (anonymous):

there we go thats what i missed than you

jimthompson5910 (jim_thompson5910):

So ANY perpendicular line to y = 4x - 3 will be of the form y = -1/4x + b where b is some constant yet to be determined

OpenStudy (anonymous):

so then the answer is -1/4x + 7/2?

jimthompson5910 (jim_thompson5910):

do you know how to determine the value of b?

OpenStudy (anonymous):

no :(

jimthompson5910 (jim_thompson5910):

y = mx + b ... start with the general linear equation y = -1/4x + b ... plug in the perpendicular slope 3 = -1/4(2) + b ... plug in x = 2, y = 3 (from (2, 3)) Now solve for b

OpenStudy (anonymous):

ok so you multiply -1/4 * 2 ?

OpenStudy (anonymous):

is D the answer?

jimthompson5910 (jim_thompson5910):

2 is the same as 2/1

jimthompson5910 (jim_thompson5910):

(-1/4)*2 = (-1/4)*(2/1)

jimthompson5910 (jim_thompson5910):

So what does (-1/4)*(2/1) turn into?

OpenStudy (anonymous):

i got 0.5

OpenStudy (anonymous):

so now what?

jimthompson5910 (jim_thompson5910):

which is what as a fraction?

OpenStudy (anonymous):

what do you mean? lol

jimthompson5910 (jim_thompson5910):

0.5 is equal to what fraction

jimthompson5910 (jim_thompson5910):

for example, 0.2 = 1/5

jimthompson5910 (jim_thompson5910):

so, 0.5 = ???

OpenStudy (anonymous):

ooh because i got that answer when i multiplied -1/4 and 2/1

jimthompson5910 (jim_thompson5910):

I know, but I want 0.5 in fraction form

jimthompson5910 (jim_thompson5910):

hint: 0.5 = five tenths = 5/10 = ???

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

not just 2

jimthompson5910 (jim_thompson5910):

10/5 = 2 but 5/10 = 1/2

jimthompson5910 (jim_thompson5910):

so 0.5 = 1/2

jimthompson5910 (jim_thompson5910):

that means 3 = -1/4(2) + b becomes 3 = -1/2 + b

OpenStudy (anonymous):

ok im getting it now so what the final answer? cause i just got confused lol

jimthompson5910 (jim_thompson5910):

well add 1/2 to both sides to isolate b

jimthompson5910 (jim_thompson5910):

What do you get when you do that?

OpenStudy (anonymous):

i think you get 5 rite/

OpenStudy (anonymous):

*?

jimthompson5910 (jim_thompson5910):

3 + 1/2 is not 5

jimthompson5910 (jim_thompson5910):

think of 3 as 3/1 Then multiply top and bottom by 2 to get 6/2

jimthompson5910 (jim_thompson5910):

3 + 1/2 3/1 + 1/2 6/2 + 1/2 ???

OpenStudy (anonymous):

7/2?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

m = -1/4 b = 7/2 so, the final answer is y = -1/4x + 7/2

OpenStudy (anonymous):

thank you